A box contains b red balls 2b white 3b blue balls $ b>0


3 balls are selected without replacement if A is the event that no two of the selected balls have same colour then


1) no value of b for which P(A)= 0.3


2) one value of b for which P(A)= 0.3 and this value is less than 10


3) two values of b and this value is greater than 10


4) None

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Mayank

3 balls are selected without replacement if  no two of the selected balls have same colour then 

 = bC1 * 2bC1 *3bC

 

Probability P =[ bC1 * 2bC1 *3bC]/6bC3

                    =6b2 /(6b-1)(6b-2)

 for P = 0.3

  6b2 /(6b-1)(6b-2)  =0.3

  8b2 -9b +1 =0

  b=1 and b =1/8

 b can not be 1/8

 so only one value of b =1

option b is correct

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin